OpenStudy (anonymous):

The volume of 0.2N feso4 required to reduce 0.01 equivalent of KMnO4 in acid medium is?

OpenStudy (anonymous):

start of with chemical reaction equation

OpenStudy (anonymous):

2Kmno4 + 8 H2so4 + 10 feso4 -> k2so4 + Mnso4 + 5fe2(so4)3 + 8h2o

OpenStudy (anonymous):

where did you get that H2SO4? it says acidic medium as i see...

OpenStudy (anonymous):

In acidic medium so..

OpenStudy (anonymous):

so write equation again without H2SO4

OpenStudy (anonymous):

Why??? we have to acid right.

OpenStudy (anonymous):

oh god, im toooo tired for this... plead to somebody else to help you

OpenStudy (anonymous):

@Preetha plz help

OpenStudy (anonymous):

I am sorry. I am so confused. I assumed you would use mass/volume formula after balancing.

OpenStudy (anonymous):

@siddhantsharan

OpenStudy (anonymous):

By law of equivilants, Equiv. of feso4 = equiv. of Kmno4. V * 0.2N = 0.01 Therefore V = 0.2/0.01 = 20litres.

OpenStudy (anonymous):

thxxx

OpenStudy (anonymous):

@Kryten told me to write the eq

OpenStudy (anonymous):

You dont need that. Thats the hard way. :P

OpenStudy (anonymous):

ok thxx a lot